### v2

@@ 1 Introduction
In the era of digital content consumption, recommender systems play an essential role in shaping user experiences and engagement on various platforms, including search engines, streaming services, social networks, and more. The core task of a recommender system is to match users with relevant content, created by content creators. By that, the system induces an ecosystem in which content providers compete for exposure and often engage in strategic behavior to maximize visibility [QJ24]. In particular, the specific recommendation mechanism used by the platform determines the incentives, and hence the strategic behavior, of the content providers. In the context of search engines, this phenomenon is often referred to as Search Engine Optimization (SEO) and is widely observed in real-world search engines [KT22].
@@
These dynamics have major implications for various economic aspects of the ecosystem. First, strategic behavior affects the welfare of both end users (seeking relevant content) and content creators, whose profits are often proportional to the level of exposure they receive from the platform. The second aspect is stability, which refers to the existence of, and the convergence to, a stable state, in which content providers have no incentive to modify their content given the content provided by their competitors. In the game-theoretic jargon, this stability is captured by the notion of Nash equilibrium. Not surprisingly, these two aspects have been shown, for example empirically in [MPRT24], to be interconnected.
@@
The study of convergence to equilibrium through the lens of game theory, a field often referred to as learning dynamics, has been adopted to analyze the stability of recommendation and ranking mechanisms. Specifically, two concrete dynamics that arise in the presence of strategic behavior over time are better-response dynamics and no-regret dynamics. The convergence of better-response dynamics has been studied in the context of information retrieval under different modeling assumptions [BPRT19, MPRT24]. In this paper, we study no-regret learning dynamics in information retrieval games. We adopt the game-theoretic framework of [YLS+24, MPRT24], in which both content and users are represented in a continuous, multi-dimensional embedding space, and the content providers aim to maximize exposure while maintaining the integrity of their original content. As in [MPRT24], and unlike in [YLS+24], we consider mechanisms that only control the exposure of each content provider, without any additional reward in the form of payments (as done in platforms such as YouTube [You23]). These mechanisms contain, among others, search engines that control the probability of presenting documents in response to a user query. That is, the mechanism considered is a ranking function which is a mapping from the publishers’ documents to a distribution over the publishers, determining each publisher’s exposure rate.
@@
In the spirit of previous work [HKJ+22, JGS23, YLS+24, MPRT24], we model SEO as a game and define a natural class of ranking functions we term proportional ranking functions. While in our model a ranking function is a vector-valued function (as it assigns exposure probabilities for each publisher), a proportional ranking function is uniquely defined by a one-variable scalar function, which we call an activation function. Importantly, the choice of the ranking function determines the incentives of the strategic publishers and hence affects the content available to users in the corpus. Relying on the notion of socially concave games, a subclass of concave games presented in [EdMN09], we prove that the concavity of the activation function guarantees the convergence of any no-regret learning dynamics in any induced SEO game. To do so, we establish a rather stronger statement by proving the equivalence of the following three conditions: the activation function is concave; any induced game is socially-concave; any induced game is concave (as in [Ros65]). We also study the empirical trade-offs between users’ and publishers’ welfare under different choices of the activation function, using a state-of-the-art no-regret dynamics algorithm [FAL+22]. In addition, we demonstrate the effect of the game parameters such as the penalty factor and the number of publishers on the publishers’ welfare, the users’ welfare and the convergence rate.
@@ 1.1 Related work
In a recent growing line of research work, a variety of game-theoretic frameworks have been proposed to analyze the strategic dynamics of content providers in recommendation systems, particularly in the context of search engines. The study of strategic behavior of content creators under the mediation of a recommender system was initiated in a series of research works [BPT17, BPT18, BPGRT19]. The Shapley mediator was proposed as a mechanism satisfying both stability (in the form of equilibrium existence guarantee) as well as fairness-related axioms.
Another concurrent line of works focused on the analysis of an adversarial information retrieval framework, in which content providers compete on exposure [BBTK15, BTK17]. These works demonstrated the sub-optimality of greedy ranking according to the probability ranking principle (PRP) of [Rob77] in terms of social welfare at equilibrium. [RRTK17] then studied both theoretically and empirically the SEO competition in a repeated setting. Under similar modeling assumptions, [BPRT19] showed that better-response dynamics of publishers in PRP-based retrieval games are guaranteed to converge.
More recent studies on strategic behavior within recommendation and search ecosystems have adapted several modeling assumptions that reflect the use of deep learning and dense representations [HKJ+22, JGS23, YLS+24, MPRT24]. Unlike earlier studies, these recent efforts assume that content is represented within a continuous embedding space, rather than a discrete set of ”topics”. In contrast to [YLS+24], we focus on the case in which the platform can only control the exposure rate of the publishers. The main focus of these works was on equilibria characterization [HKJ+22, JGS23] and the study of better-response dynamics convergence [MPRT24] under specific recommendation or ranking mechanisms. In contrast, this work characterizes a rich and intuitive family of mechanisms that guarantee the convergence of no-regret dynamics.
The fundamental connections between no-regret learning and game-theoretic solution concepts [FV97, FS99, HMC00, BHLR08, Rou15] have made the regret-minimization framework very appealing to the machine learning and artificial intelligence communities [RS13, BBJT15, SALS15, MSB+17, BS18, KHSC18, DFG21, FAL+22, APFS24]. In particular, the notion of concave games [Ros65] and its connection to the convergence of no-regret dynamics showed by [EdMN09] plays a crucial role in our analysis.
@@ 2 Preliminaries
No-regret dynamics is a well-known type of dynamics within the online learning framework, in which each player $i$ is a learning agent that has to select a strategy $x_{i}^{(t)}\in X_{i}$ in every round $t\in\operatorname*{\mathbb{N}}$, in response to the sequence of the observations she has received up until round $t$. A widely acceptable way to measure the performance of player $i$ up to some horizon time $T\in\operatorname*{\mathbb{N}}$ in this setting is regret, [^1] We consider the notion of external regret, not to be confused with internal (’swap’) regret. defined as follows:
@@
That is, the regret of player $i$ is measured with respect to the utility she could have gained had she selected in every round the best fixed strategy in hindsight. In our context, as standard in the online learning literature, e.g. [FAL+22], it is assumed that each player observes in every round $t$ the gradient of her utility function with respect to her strategy, evaluated at the strategy profile selected by all players in this round. That is, in round $t$ player $i$ observes $\nabla_{x_{i}}u_{i}(x^{(t)})$.
@@
We now introduce two special classes of games. A game $G$ is said to be a (strictly) concave game if for any player $i$, $u_{i}(x_{i},x_{-i})$ is (strictly) concave in $x_{i}$ for any fixed $x_{-i}$. In his seminal work, [Ros65] introduced the class of concave games, and showed that every (strictly) concave game has a (unique) Nash equilibrium. [^2] Concave games are well studied, and possess some other nice properties, see e.g. [Ui08, EHL22]. An important subclass of concave games is the class of socially-concave games, introduced by [EdMN09].
@@
    There exist strictly positive coefficients $\alpha_{1},...\alpha_{n}>0$, such that $\sum_{i=1}^{n}\alpha_{i}=1$ and the function $g(x)\coloneqq\sum_{i=1}^{n}\alpha_{i}\cdot u_{i}(x)$ is concave in $x$.
@@
[EdMN09] showed that every socially-concave game is a concave game, but the reverse direction is not necessarily true. [EdMN09] also showed that if each player follows any regret minimization procedure in a socially-concave game, then the dynamics are guaranteed to converge. The convergence is in the sense that both the average action vector converges to a Nash equilibrium and the average utility of each player converges to her utility in that Nash equilibrium.
@@ 3 The model
Information needs of the users are denoted by $x^{*}\in[0,1]^{k}$ and modeled as vectors in the same embedding space as the publishers’ documents. This approach mirrors the common practice in standard models of dense retrieval [ZML+20, ZLRW22]. We model the fact that in real life there are often multiple information needs by a demand distribution $P^{*}\in\mathcal{P}([0,1]^{k})$, a distribution of information needs. In addition, each publisher has an initial document $x_{0}^{i}\in[0,1]^{k}$.
Distance between two documents in the embedding space is measured by a given semi-metric [^3] A semi-metric is similar to a metric, except it does not need to satisfy the triangle inequality. $d:[0,1]^{k}\times[0,1]^{k}\to\operatorname*{\mathbb{R}}_{+}$, which is assumed to be normalized such that for any two documents $a,b\in[0,1]^{k}$ we have $0\leq d(a,b)\leq 1$. Another standard assumption that we make is that $d(a,x_{0}^{i})$ and $d(a,x^{*})$ are twice differentiable in $a$ for all the initial documents $\{x_{0}^{i}\}_{i\in N}$ and for every information need $x^{*}$. This assumption is needed for the utilities to be twice differentiable. Moreover, we assume that $d(a,b)$ is bi-convex, meaning that it is convex in $a$ for any fixed $b$ and vice-versa. The role of $d$ in our model is two-fold. Firstly, it measures the integrity to the initial content, with the interpretation that a small distance between a publisher’s strategy document and her initial document represents strong integrity. Secondly, $d$ is used by the system designer to assess the relevance of documents to a given information need $x^{*}$, with the interpretation that $d$-proximity implies relevance.
@@
Note that condition A1. of social concavity holds with $\alpha_{i}=\frac{1}{n}$ in every publishers’ game, since $\sum_{i=1}^{n}r_{i}\equiv 1$. This leads to the following useful lemma, whose proof is deferred to Appendix A.1.
@@
We say a ranking function $r$ is a proportional ranking function if there exists a twice differentiable and strictly decreasing function $g:[0,1]\to\operatorname*{\mathbb{R}}_{++}$ such that
@@
Our main objective is to characterize the subset of proportional ranking functions that induce socially-concave games. These ranking functions possess various desirable properties, such as the existence of a Nash equilibrium, and a guarantee that if publishers engage according to any regret-minimization procedure, they will converge to a Nash equilibrium. We can then simulate no-regret dynamics for various instances of the publishers’ game for different ranking functions, and compare them in terms of publishers’ and users’ welfare.
@@ 4 Main result
We now provide a necessary and sufficient condition for a proportional ranking function to induce socially-concave games. This condition turns out to have a relatively simple form: the activation function should be concave. We show this by proving a stronger statement:
@@
Let $r$ be a proportional ranking function with activation function $g$. Then, the following are equivalent:
@@
[…]
$r_{i}(x_{i},x_{-i};x^{*})=\frac{g\bigl{(}d(x_{i},x^{*})\bigr{)}}{\sum_{j=1}^{n}g\bigl{(}d(x_{j},x^{*})\bigr{)}}=\frac{C}{C+h(x_{-i})}$
$r_{i}(x_{i},x_{-i};x^{*})=\ell\bigl{(}h(x_{-i})\bigr{)}$
[…]
Theorem 1, together with the findings of [EdMN09] that in socially-concave games any no-regret dynamics converges, and the guarantee of [Ros65] that (strictly) concave games possess a (unique) NE, leads to the following result:
@@
Let $r$ be a proportional ranking function with activation function $g$. Then, if $g$ is concave, any no-regret learning dynamics in any publishers’ game induced by $r$ converges. Moreover, any induced publishers’ game possesses a NE, and if $d$ is strictly bi-convex the NE is unique.

### v3

@@ 1 Introduction
In the era of digital content consumption, recommender systems play an essential role in shaping user experiences and engagement on various platforms, including search engines, streaming services, social networks, and more. The core task of a recommender system is to match users with relevant content, created by content creators. By that, the system induces an ecosystem in which content providers compete for exposure and often engage in strategic behavior to maximize visibility (Qian & Jain, 2024). In particular, the specific recommendation mechanism used by the platform determines the incentives, and hence the strategic behavior, of the content providers. In the context of search engines, this phenomenon is often referred to as Search Engine Optimization (SEO) and is widely observed in real-world search engines (Kurland & Tennenholtz, 2022).
@@
These dynamics have major implications for various economic aspects of the ecosystem. First, strategic behavior affects the welfare of both end users (seeking relevant content) and content creators, whose profits are often proportional to the level of exposure they receive from the platform. The second aspect is stability, which refers to the existence of, and the convergence to, a stable state, in which content providers have no incentive to modify their content given the content provided by their competitors. In the game-theoretic jargon, this stability is captured by the notion of Nash equilibrium. Not surprisingly, these two aspects have been shown to be interconnected in Madmon et al. (2023).
The study of convergence to equilibrium through the lens of game theory, a field often referred to as learning dynamics, has been adopted to analyze the stability of recommendation and ranking mechanisms. Specifically, two concrete dynamics that arise in the presence of strategic behavior over time are better-response dynamics and no-regret dynamics. The convergence of better-response dynamics has been studied in the context of information retrieval under different modeling assumptions (Ben-Porat et al., 2019b; Madmon et al., 2023).
In this paper, we study no-regret learning dynamics in information retrieval games, as they are particularly relevant for modeling SEO behavior. In many practical applications, publishers use the service of SEO experts, who employ advanced techniques to maximize the visibility of their content. These experts are judged based on hindsight—specifically, how much more they could have achieved had they known their competitors’ strategies in advance. This retrospective, backward-looking evaluation, which SEO strategies aim to minimize, is precisely encapsulated by the concept of regret.
@@
In the spirit of previous work (Hron et al., 2022; Jagadeesan et al., 2023; Madmon et al., 2023; Yao et al., 2023; 2024a; 2024b; 2024c), we take a game-theoretic approach and model SEO as a game. In our model, both content and users are represented in a continuous, multi-dimensional embedding space, and the content providers aim to maximize exposure while maintaining the integrity of their original content. As in Madmon et al. (2023), we consider mechanisms that only control the exposure of each content provider, without any additional reward in the form of payments (as done in platforms such as YouTube, 2023). These mechanisms contain, among others, search engines that control the probability of presenting documents in response to a user query. That is, the mechanism considered is a ranking function which is a mapping from the publishers’ documents to a distribution over the publishers, determining each publisher’s exposure rate.
@@
We define a natural class of ranking functions we term proportional ranking functions (PRFs). While in our model a ranking function is a vector-valued function (as it assigns exposure probabilities for each publisher), a proportional ranking function is uniquely defined by a one-variable scalar function, which we call an activation function. Importantly, the choice of the ranking function determines the incentives of the strategic publishers and hence affects the content available to users in the corpus.
Relying on the notion of socially concave games, a subclass of concave games presented in Even-dar et al. (2009), we prove that the concavity of the activation function guarantees the convergence of any no-regret learning dynamics in any induced SEO game. To do so, we establish a rather stronger statement by proving the equivalence of the following three conditions: the activation function is concave; any induced game is socially concave; any induced game is concave (as in Rosen, 1965). This result can also be interpreted as addressing the learnability of equilibrium strategies’ embedding representations from the publishers’ perspective. It highlights how platform designers can devise ranking schemes that induce ecosystems in which publishers can learn an equilibrium through no-regret dynamics. This learnability result also serves as a stability guarantee, since the publishers know that they will not gain much by deviating from the learned equilibrium. Importantly, as the learned equilibrium is pure, this means the publishers will play deterministically, thus achieving stability.
We also study the empirical trade-offs between users’ and publishers’ welfare under different choices of the activation function, using a state-of-the-art no-regret dynamics algorithm (Farina et al., 2022). In addition, we demonstrate the effect of the game parameters such as the penalty factor and the number of publishers on the publishers’ welfare, the users’ welfare and the convergence rate.
@@ 1.1 Related work
In a recent growing line of research work, a variety of game-theoretic frameworks have been proposed to analyze the strategic dynamics of content providers in recommendation systems, particularly in the context of search engines. The study of strategic behavior of content creators under the mediation of a recommender system was initiated in a series of research works (Ben-Porat & Tennenholtz, 2017; 2018; Ben-Porat et al., 2019a). The Shapley mediator was proposed as a mechanism satisfying both stability (in the form of equilibrium existence guarantee) as well as fairness-related axioms.
Another concurrent line of works focused on the analysis of an adversarial information retrieval framework, in which content providers compete on exposure (Ben Basat et al., 2015; 2017). These works demonstrated the sub-optimality of greedy ranking according to the probability ranking principle (PRP) of Robertson (1977) in terms of social welfare at equilibrium. Raifer et al. (2017) then studied both theoretically and empirically the SEO competition in a repeated setting. Under similar modeling assumptions, Ben-Porat et al. (2019b) showed that better-response dynamics of publishers in PRP-based retrieval games are guaranteed to converge.
More recent studies on strategic behavior within recommendation and search ecosystems have adapted several modeling assumptions that reflect the use of deep learning and dense representations (Hron et al., 2022; Jagadeesan et al., 2023; Madmon et al., 2023; Yao et al., 2023; 2024a; 2024b; 2024c). Unlike earlier studies, these recent efforts assume that content is represented within a continuous embedding space, rather than a discrete set of "topics". In contrast to Yao et al. (2024b), we focus on the case in which the platform can only control the exposure rate of the publishers. The main focus of these works was on equilibria characterization (Hron et al., 2022; Jagadeesan et al., 2023; Yao et al., 2024a; b; c) and the study of learning dynamics (Madmon et al., 2023; Yao et al., 2023) under specific exposure or rewarding mechanisms. In contrast, this work characterizes a general, rich and intuitive family of mechanisms that guarantee the convergence of no-regret dynamics to a Nash equilibrium. [^1] While Yao et al. (2023) also studied no-regret dynamics, their analysis is focused on a particular structure of publishers’ incentives. They also did not study stability criteria such as convergence to a Nash equilibrium, but instead focused on studying average welfare over time.
The fundamental connections between no-regret learning and game-theoretic solution concepts (Foster & Vohra, 1997; Freund & Schapire, 1999; Hart & Mas-Colell, 2000; Blum et al., 2008; Roughgarden, 2015) have made the regret-minimization framework very appealing to the machine learning and artificial intelligence communities (Rakhlin & Sridharan, 2013; Bowling et al., 2015; Syrgkanis et al., 2015; Moravčík et al., 2017; Brown & Sandholm, 2018; Kangarshahi et al., 2018; Daskalakis et al., 2021; Farina et al., 2022; Anagnostides et al., 2024). In particular, the notion of concave games (Rosen, 1965) and its connection to the convergence of no-regret dynamics showed by Even-dar et al. (2009) plays a crucial role in our analysis.
@@ 2 Preliminaries
No-regret dynamics is a well-known type of dynamics within the online learning framework, in which each player $i$ is a learning agent that has to select a strategy $x_{i}^{(t)}\in X_{i}$ in every round $t\in\operatorname*{\mathbb{N}}$, in response to the sequence of the observations she has received up until round $t$. A widely accepted way to measure the performance of player $i$ up to some horizon time $T\in\operatorname*{\mathbb{N}}$ in this setting is regret, [^2] We consider the notion of external regret, not to be confused with internal (’swap’) regret. defined as follows:
@@
That is, the regret of player $i$ is measured with respect to the utility she could have gained had she selected in every round the best fixed strategy in hindsight. In our context, as standard in the online learning literature, e.g. Farina et al. (2022), it is assumed that each player observes in every round $t$ the gradient of her utility function with respect to her strategy, evaluated at the strategy profile selected by all players in this round. That is, in round $t$ player $i$ observes $\nabla_{x_{i}}u_{i}(x^{(t)})$. A no-regret dynamics is a dynamics in which for any player $i$, $\text{Reg}_{i}^{T}=o(T)$, meaning that the player’s regret grows sub-linearly with respect to the time horizon $T$.
@@
We now introduce two special classes of games. A game $G$ is said to be a (strictly) concave game if for any player $i$, $u_{i}(x_{i},x_{-i})$ is (strictly) concave in $x_{i}$ for any fixed $x_{-i}$. In his seminal work, Rosen (1965) introduced the class of concave games, and showed that every (strictly) concave game has a (unique) Nash equilibrium. [^3] Concave games are well studied, and possess some other properties; see, e.g., Ui (2008); Einy et al. (2022). An important subclass of concave games is the class of socially-concave games, introduced by Even-dar et al. (2009).
@@
    There exist strictly positive coefficients $\alpha_{1},\ldots,\alpha_{n}>0$, such that $\sum_{i=1}^{n}\alpha_{i}=1$ and the function $g(x)\coloneqq\sum_{i=1}^{n}\alpha_{i}\cdot u_{i}(x)$ is concave in $x$.
@@
Even-dar et al. (2009) showed that every socially-concave game is a concave game, but the reverse direction is not necessarily true. Even-dar et al. (2009) also showed that if each player follows any regret minimization procedure in a socially-concave game, then the dynamics are guaranteed to converge. The convergence is in the sense that the average action vector converges to a Nash equilibrium.
Theorem 1.
(Even-dar et al., 2009) Let $\{x^{(t)}\}_{t=1}^{T}$ be a sequence of strategy profiles in a socially-concave game. Then the average strategy profile vector $\hat{x}^{(T)}\coloneqq\frac{1}{T}\sum_{t=1}^{T}x^{(t)}$ is an $\varepsilon_{T}$-NE, where $\varepsilon_{T}\coloneqq\frac{1}{\alpha_{\text{min}}}\sum_{j\in N}\frac{\alpha_{j}\text{Reg}_{j}^{T}}{T}$ and $\alpha_{\text{min}}\coloneqq\min_{j\in N}\alpha_{j}$.
@@ 3 The model
Information needs of the users are denoted by $x^{*}\in[0,1]^{k}$ and modeled as vectors in the same embedding space as the publishers’ documents. This approach mirrors the common practice in standard models of dense retrieval (Zhan et al., 2020; Zhao et al., 2022). We model the fact that in real life there are often multiple information needs by a demand distribution $P^{*}\in\mathcal{P}([0,1]^{k})$, a distribution of information needs. In addition, each publisher has an initial document $x_{0}^{i}\in[0,1]^{k}$.
Distance between documents in the embedding space is measured by a given semi-metric [^4] A semi-metric is similar to a metric, except it does not need to satisfy the triangle inequality. $d:[0,1]^{k}\times[0,1]^{k}\to\operatorname*{\mathbb{R}}_{+}$, which is assumed to be normalized such that $0\leq d(a,b)\leq 1$ for any two documents $a,b\in[0,1]^{k}$. Another standard assumption that we make is that $d(a,x_{0}^{i})$ and $d(a,x^{*})$ are twice differentiable in $a$ for all the initial documents $\{x_{0}^{i}\}_{i\in N}$ and for every information need $x^{*}$ in the support of $P^{*}$. This assumption is needed for the utilities to be twice differentiable. Moreover, we assume that $d(a,b)$ is bi-convex, meaning that it is convex in $a$ for any fixed $b$ and vice-versa. The role of $d$ in our model is two-fold. Firstly, it measures the integrity to the initial content, with the interpretation that a small distance between a publisher’s strategy document and her initial document represents strong integrity. Secondly, $d$ is used by the system designer to assess the relevance of documents to a given information need $x^{*}$, with the interpretation that $d$-proximity implies relevance.
@@
Note that condition A1. of social concavity holds with $\alpha_{i}=\frac{1}{n}$ in every publishers’ game, since $\sum_{i=1}^{n}r_{i}\equiv 1$. This leads to the following useful lemma, whose proof is deferred to Appendix B.1.
@@
We say a ranking function $r$ is a proportional ranking function (PRF) if there exists a twice differentiable and strictly decreasing function $g:[0,1]\to\operatorname*{\mathbb{R}}_{++}$ such that
@@
Our main objective is to characterize the subset of PRFs that induce socially-concave games. These ranking functions possess various desirable properties, such as the existence of a Nash equilibrium, and a guarantee that if publishers engage according to any regret-minimization procedure, they will converge to a Nash equilibrium. We can then simulate no-regret dynamics for various instances of the publishers’ game for different ranking functions, and compare them in terms of publishers’ and users’ welfare.
@@ 4 Main result
We now provide a necessary and sufficient condition for a PRF to induce socially-concave games. This condition turns out to have a relatively simple form: the activation function should be concave. We show this by proving a stronger statement:
@@
Let $r$ be a PRF with activation function $g$. Then, the following are equivalent:
@@
Theorem 2, together with the findings of Even-dar et al. (2009) that in socially-concave games any no-regret dynamics converges, and the guarantee of Rosen (1965) that (strictly) concave games possess a (unique) NE, leads to the following result:
@@
Let $r$ be a PRF with activation function $g$. Then, if $g$ is concave, any no-regret learning dynamics in any publishers’ game induced by $r$ converges. Moreover, any induced publishers’ game possesses an NE, and if $d$ is strictly bi-convex the NE is unique. [^7] In Appendix D we characterize this NE and provide a preliminary welfare analysis.
The strict bi-convexity assumption on the semi-metric is satisfied, for instance, by the squared Euclidean norm, which is the semi-metric we use in the simulations conducted in §5. Corollary 1 implies that the system designer can readily guarantee convergence in average of no-regret dynamics, by simply choosing a concave activation function. While in §5 we empirically show that the selection of the specific activation function has a significant effect on various evaluation criteria, convergence of the average profile to an $\varepsilon$-NE is guaranteed for any concave activation.
Equilibrium strategy learning
We now discuss a key practical implication of our main result for real-world platforms, such as recommendation systems and search engines. As established, ranking with concave activation only ensures that the average document profile converges to equilibrium, which is a weaker notion of convergence compared to the convergence of the final iteration’s strategy profile. This may seem counter to the platform’s goal of stability, where reducing fluctuations over time in publishers’ content is a priority. However, from a practical standpoint, our main result can be seen as a positive result regarding the learnability of an approximate NE from the publishers’ perspective. Assume that each publisher $i$ has a budget that allows her to hire an SEO expert, who utilizes a no-regret algorithm, for $T_{i}$ time periods (determined, for example, by budget constraints). In Proposition 1 we establish that, under the assumption that the stopping times are not too far from each other, the publishers can learn an approximate NE.
Proposition 1.
Let $G$ be a socially-concave publishers’ game. Suppose each player $i$ accumulates $\text{Reg}_{i}^{T_{i}}$ regret in the first $T_{i}$ rounds. Let $x^{\text{eq}}_{i}\coloneqq\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}x^{(t)}_{i}$. Then $x^{\text{eq}}=\left(x^{\text{eq}}_{1},\ldots,x^{\text{eq}}_{n}\right)$ is an $\varepsilon$-NE with $\varepsilon=\frac{1}{T_{\text{max}}}\sum_{j\in N}\left(\text{Reg}_{j}^{T_{j}}+\left(1+\lambda_{j}\right)\left(T_{\text{max}}-T_{j}\right)\right)$, where $T_{\text{max}}=\max_{i\in N}T_{i}$.
This result, whose proof is deferred to Appendix B.3, implies that given that the stopping times are sufficiently large, and not too far apart, any publisher can efficiently learn an equilibrium strategy. This means that each publisher knows that she can play deterministically the strategy she learned, without expecting significant gains from deviating to any other strategy. Consequently, content creation becomes deterministic in the long run, eliminating fluctuations within the corpus and achieving stability. [^8] This also explains why convergence to a correlated equilibrium (CE) may be insufficient for stability. While it is well-known that no-regret dynamics converge to CE in general games (in the same ”average” sense), such equilibria allow for non-deterministic actions, leading to the same issue of frequent content modifications over time. Such a stability guarantee ensures predictability regarding the content available in the corpus for both the users and the ecosystem designer, which is crucial for the design of reliable search and recommendation ecosystems.